OpenStudy (anonymous):

Can someone help me with this? "A 2.0-kg cart and a 3.0-kg cart collide on a low-friction track. The 3.0-kg cart is initially moving at 1.5m/s to the right, but after the collision it is moving at 5.5m/s to the right. After the collision, the 2.0-kg cart is moving to the right at 3.0m/s. If the direction from left to right is positive, what was the 2.0-kg cart's initial velocity?"

OpenStudy (zpupster):

i believe this is an inelastic collision. Since the two cars stick together, both are moving to the right--if i am thinking about this correctly it is a rear end collision A into B using the conservation of momentum is enough to solve for our one unknown variable- m1v1 + m2v2 = Mvf so M = added masses (2.0)(v1) + (3.0)(1.5) = (5.0)(5.5) v1=??